# Layered Insights: Generalizable Analysis of Human Authorial Style by Leveraging All Transformer Layers

### v2

@@ Abstract
We propose a new approach for the authorship attribution task that leverages the various linguistic representations learned at different layers of pre-trained transformer-based models. We evaluate our approach on two popular authorship attribution models and three evaluation datasets, in in-domain and out-of-domain scenarios. We found that utilizing various transformer layers improves the robustness of authorship attribution models when tested on out-of-domain data, resulting in new state-of-the-art results. Our analysis gives further insights into how our model’s different layers get specialized in representing certain stylistic features that benefit the model when tested out of the domain.
@@ 1 Introduction
To evaluate our hypothesis, we follow the same evaluation setup of Rivera-Soto et al. (2021) that considers three datasets for evaluation: Reddit comments, Amazon reviews, and Fanfiction short stories. We apply our approach to two popular AA models: LUAR Rivera-Soto et al. (2021) and Wegmann Wegmann et al. (2022), considering two base transformers for each: RoBERTa Liu (2019) and ModernBERT Warner et al. (2024). Our experiments demonstrate that, for the out-of-domain scenario, applying our approach on two AA models with two base transformers leads to improvement in 15 out of 16 scenarios, with the highest improvement of around 50% in the mean reciprocal rank (MRR) for LUAR when trained on the Fanfiction dataset and evaluated on Reddit. Results, however, in the in-domain scenario are mixed. These insights indicate strong generalization capabilities gained when applying our approach, which we argue is more relevant in practical scenarios of the AA task than in-domain. This also provides evidence of the importance of leveraging the various layers of the base transformer when working on the AA task. Furthermore, we perform experiments to analyze the role each layer plays in the overall performance of our model. We found that certain layers became more relevant when models are tested in the out-of-domain setting, leading to high performance. The code for our experiments will be publicly available upon acceptance.
In summary, our contributions are the following:
@@ 6 Analysis
In this section, we look at what our model’s projection layers capture linguistically and what role they play in the final prediction in both in and out-of-domain scenarios. All analyses are run when our approach is applied to the LUAR (LIGHT-LUAR) model.
@@ 6.1 Learned Linguistic Representations
Figure 2 shows a heatmap of the linguistic competence of all layers of the two models. We can observe that LUAR’s output layer (first column) has a huge skew towards morphology compared to other linguistic phenomena, indicating the importance of morphological representation in texts for the AA task. In contrast, our LIGHT-LUAR model has a more balanced representation of all the linguistic phenomena, with the highest being morphology captured at Layer 4 (L 4). Aside from Morphology, our LIGHT-LUAR represents each linguistic phenomenon at one of its layers better than LUAR’s final output layer. Aligned with previous literature, Syntax is best captured at the intermediate layers (L 2-3), while discourse is at the later layers (L 3-6), as well as L 0, potentially, capturing discourse markers. These observations indicate the need to model authorship attribution at different layers to benefit from the diverse captured linguistic representation, rather than considering only the final output layer.
@@ 7 Discussion and Conclusion
We evaluated our approach by applying it to two popular AA models and evaluated it on three datasets. Results confirmed how our approach leads to increased performance on the task in the out-of-domain scenario while maintaining relatively strong performance in the in-domain settings. We then studied what linguistic representations were captured at the different layers of our model, demonstrating the reason behind its reliable and robust performance. We further analyze the contribution of each layer in our model towards the final prediction through an ablation study. Across the three trained models, we found a trend of having layers contributing more to the out-of-domain performance than in-domain. Later layers of all models were found to be essential to the in-domain performance.
@@ 8 Limitations
First, our model architecture implies adding extra parameters to the original AA structure that require more training time and computational resources, especially when the number of layers of the base transformer increases. Future experiments could consider whether all layers need to be projected in order to achieve the best performance. Second, although we verified our hypothesis on two different AA models with two different base transformers, we focused the presentation of our analysis only on the LUAR model due to space limitations. Third, we rely on probing methods to analyze the linguistic competence of our projected layers. However, these might not be very reliable tools, and they have their own limitations.

### v3

@@ Abstract
We propose a new approach for the authorship attribution task that leverages the various linguistic representations learned at different layers of pre-trained transformer-based models. We evaluate our approach on two popular authorship attribution models and three evaluation datasets, in in-domain and out-of-domain scenarios. We found that utilizing various transformer layers improves the robustness of authorship attribution models when tested on out-of-domain data, resulting in a much stronger performance. Our analysis gives further insights into how our model’s different layers get specialized in representing certain linguistic aspects that we believe benefit the model when tested out of the domain.
**footnotetext: These authors contributed equally to this work§§footnotetext: Author is currently working at JPMorgan Chase & Co.
@@ 1 Introduction
To evaluate our hypothesis, we follow the same evaluation setup of Rivera-Soto et al. (2021) that considers three datasets for evaluation: Reddit comments, Amazon reviews, and Fanfiction short stories. We apply our approach to two popular AA models: LUAR Rivera-Soto et al. (2021) and Wegmann Wegmann et al. (2022), considering two base transformers for each: RoBERTa Liu (2019) and ModernBERT Warner et al. (2024). Our experiments demonstrate that, for the out-of-domain scenario, applying our approach on two AA models with two base transformers leads to improvement in 15 out of 16 scenarios, with the highest improvement of around 50% in the mean reciprocal rank (MRR) for LUAR when trained on the Fanfiction dataset and evaluated on Reddit. Results, however, in the in-domain scenario are mixed. These insights indicate strong generalization capabilities gained when applying our approach, which we argue is more relevant in practical scenarios of the AA task than in-domain. This also provides evidence of the importance of leveraging the various layers of the base transformer when working on the AA task. Furthermore, we perform experiments to analyze the role each layer plays in the overall performance of our model. We found that certain layers became more relevant when models are tested in the out-of-domain setting, leading to high performance. In summary, our contributions are: [^1] https://github.com/MiladAlshomary/LIGHT-for-AA-analysis
@@ 5 Results
Figure 2: The linguistic capability of LUAR model compared to the best performing layers of our LIGHT-LUAR model trained on the Reddit dataset.
@@ 6 Analysis
In this section, we look at what our model’s projection layers capture linguistically and what role they play in the final prediction in both in and out-of-domain scenarios. All analyses are performed on our approach when applied to LUAR.
@@ 6.1 Learned Linguistic Representations
Figure 2 shows the best performance of LUAR in representing each of the linguistic phenomena compared to the best performing layers of LIGHT-LUAR. We can observe that LUAR’s output layer has a huge skew towards morphology compared to other linguistic phenomena, indicating the importance of morphological representation in texts for the AA task. In contrast, our LIGHT-LUAR model has a more balanced representation of all the linguistic phenomena. More details are shown in Figure 3 as a heatmap of the linguistic competence of all layers of the two models. Aside from Morphology, our LIGHT-LUAR represents each linguistic phenomenon at one of its layers better than LUAR’s final output layer. Aligned with previous literature, Syntax is best captured at the intermediate layers (L 2-3), while discourse is at the later layers (L 3-6), as well as L 0, potentially, capturing discourse markers. These observations indicate the need to model authorship attribution at different layers to benefit from the diverse captured linguistic representation, rather than considering only the final output layer.
@@ 7 Discussion and Conclusion
We evaluated our approach by applying it to two popular AA models and evaluated it on three datasets. Results confirmed how our approach leads to increased performance on the task in the out-of-domain scenario while maintaining relatively strong performance in the in-domain settings. We then studied what linguistic representations were captured at the different layers of our model, demonstrating the reason behind its reliable and robust performance. We further analyze the contribution of each layer in our model towards the final prediction through an ablation study. Across the three trained models, we found a trend of having layers contributing more to the out-of-domain performance than in-domain. Later layers of all models were found to be essential to the in-domain performance. Overall, our findings demonstrate the importance of leveraging all transformer layers when addressing the authorship attribution task to achieve robust performance in out-of-domain.
It is worth drawing parallels between our approach and the classical stylometry approaches for the authorship attribution task Stamatatos (2009). While the latter explicitly models the linguistic clues in texts, our approach does this implicitly, where each layer might attend to a different linguistic phenomenon. Future work can evaluate our approach’s behavior against these models to have better insights into how much they align in terms of performance.
@@ 8 Limitations
Generally, our study focused on the English language, and our findings might not be transferable to other languages. Further research can explore this potential. Moreover, our model architecture implies adding extra parameters to the original AA structure that require more training time and computational resources, especially when the number of layers of the base transformer increases. Future experiments could consider whether all layers need to be projected in order to achieve the best performance. Second, although we verified our hypothesis on two different AA models with two different base transformers, we focused the presentation of our analysis only on the LUAR model due to space limitations. Third, we rely on probing methods to analyze the linguistic competence of our projected layers. However, these might not be very reliable tools, and they have their own limitations.
@@
Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein